# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Appendix I Proof of Lemma 11
where the support of $\Sigma^{\nu}_{1}$ is entirely to the right of $\sigma^{\nu}_{i}$, and thus constitutes an equivalent string order parameter $\mathsf{S}_{\nu,t}$ (N.173) of the same form as Eq. 15 with a new left endpoint, where the left-endpoint operators for $\nu=x,z$ are guaranteed to anticommute by Eq. N.172.
Importantly, we observe that the new string-order parameter in Eq. N.173 has the same important properties as the naïve counterpart (15). Both objects have (i) unit expectation value in the resource state $\lvert\Psi_{t}\rangle$ (36); (ii) a right-endpoint operator given by $\sigma^{\nu}_{f}$; (iii) a left-endpoint operator of the form $\mathcal{U}\,\tau^{\nu}\,\mathcal{U}^{\dagger}$, where the objects $\tau^{\nu}$ satisfy the Pauli algebra; and (iv) all of the operators between the two endpoint commute for different values of $\nu$.
There are two key features that facilitate iterating this procedure to recover new endpoints. The first is that the left-endpoint operators (N.174) are of the same form as the original Pauli operator in Eq. 15. The second is that there exists a “leftmost” choice of $\Sigma^{\nu}_{1}$ (N.172) such that the clustered observables $\bar{A}_{\nu,3}$ through $\bar{A}_{\nu,m}$ are not connected to $\Sigma^{\nu}_{1}$ under $\mathcal{U}$. Essentially, a set of equivalent operators recovers upon multiplying $\Sigma^{\nu}_{1}$ (N.172) by stabilizer elements $\mathsf{S}_{0}$ of $\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$; the “leftmost” left-endpoint operator (N.174) is the operator $\Sigma^{\nu}_{1}\,\mathsf{S}_{0}$ for $\mathsf{S}_{0}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$ whose support is closest to—but disjoint from—the initial site $i$ (i.e., whose right weight is closest to $i$).
@@
$\Sigma^{\nu}_{1}\,\Sigma^{\nu}_{2}=\mathcal{U}^{\dagger}\,\bar{A}_{\nu,2}\,\mathcal{U}\,,~{}~{}$ (N.175)
@@
$\sigma^{\nu}_{i}\,\Sigma^{\nu}_{s}=\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\cdots\bar{A}_{\nu,s}\,\mathcal{U}\,,~{}~{}$ (N.176)
@@
$\Sigma^{\nu}_{s-1}\,\Sigma^{\nu}_{s}=\mathcal{U}^{\dagger}\,\bar{A}_{\nu,s}\,\mathcal{U}\,,~{}$ (N.177)
@@
One can iterate this procedure until all bulk operators except $\bar{A}_{\nu,m}$ have been exhausted. Choosing the leftmost form of each $\Sigma^{\nu}_{s}$ guarantees that each $\Sigma^{\nu}_{s}$ only has support to the right of $\Sigma^{\nu}_{s-1}$, and thus constitutes a genuinely new endpoint. This choice only affects the total number of measurement “regions” $m$, which delineate endpoints.
@@
$\mathsf{S}_{\nu,t,s}$ $=\left(\mathcal{U}\,\Sigma^{\nu}_{s}\,\mathcal{U}^{\dagger}\right)\,\bar{A}_{\nu,s+1}\cdots\bar{A}_{\nu,m}\,\left(\sigma^{\nu}_{f}\right)$ (N.178)
for all $s\in[1,m]$, such that $\left\langle\Psi_{t}\middle|\mathsf{S}_{\nu,t,s}\middle|\Psi_{t}\right\rangle=1$ (60) indicates perfect string order of $\lvert\Psi_{t}\rangle$ (36).
We next identify new right endpoints. Returning to Eq. N.171, using the same operators $\Sigma^{\nu}_{s}$ defined in Eqs. N.176 and N.177. Importantly, we take $\Sigma^{\nu}_{s}$ to be the operator $\Sigma^{\nu}_{s}\,\mathsf{S}_{0}$ for $\mathsf{S}_{0}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$ with the leftmost right weight (i.e., whose support is farthest to the left). All such operators are equivalent, since these quantities are (effectively) evaluated in the initial state $\lvert\Psi_{0}\rangle$ (6), so $\mathsf{S}_{0}\cong\mathds{1}$.
@@
$\mathsf{S}^{\prime}_{\nu,t,s}$ $\equiv\mathsf{S}_{\nu,t,0}\,\mathsf{S}^{\dagger}_{\nu,t,s}$
$=\left(\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}\right)$
×$\times\left(\sigma^{\nu}_{f}\,\bar{A}_{\nu,m}\cdots\bar{A}_{\nu,s+1}\,\mathcal{U}\,\Sigma^{\nu}_{s}\,\mathcal{U}^{\dagger}\right)$
$=\left(\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\right)\,\bar{A}_{\nu,1}\cdots\bar{A}_{\nu,s}\,\left(\mathcal{U}\,\Sigma^{\nu}_{s}\,\mathcal{U}^{\dagger}\right)\,\,,~{}~{}$ (N.179)
which has the same left-endpoint operator $\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}$ as $\mathsf{S}_{\nu,t,0}$ (15) but distinct right-endpoint operators $\Sigma^{\nu}_{s}$, which are the same endpoint operators in Eq. 57.
@@
$\mathcal{S}_{\nu,a,b}$ $\equiv\mathsf{S}_{\nu,t,a}\,\mathsf{S}^{\dagger}_{\nu,t,b}$ (N.180)
$=\left(\mathcal{U}\,\Sigma^{\nu}_{a}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu,a+1}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}\right)$ (N.181)
×$\times\left(\sigma^{\nu}_{f}\,\bar{A}_{\nu,m}\cdots\bar{A}_{\nu,b+1}\,\mathcal{U}\,\Sigma^{\nu}_{b}\,\mathcal{U}^{\dagger}\right)$ (N.182)
$=\left(\mathcal{U}\,\Sigma^{\nu}_{a}\,\mathcal{U}^{\dagger}\right)\,\bar{A}_{\nu,a+1}\cdots\bar{A}_{\nu,b}\,\left(\mathcal{U}\,\Sigma^{\nu}_{b}\,\mathcal{U}^{\dagger}\right)\,\,,~{}~{}$ (57)
and since the support of $\Sigma^{\nu}_{s+1}$ is entirely to the right of $\Sigma^{\nu}_{s}$, the support of $\Sigma^{\nu}_{s}$ is to the left of $\Sigma^{\nu}_{s+1}$. Hence, the operators $\Sigma^{\nu}_{s}$ constitute distinct left and right endpoint operators for each $s\in[1,m]$, where *$\left\langle*\middle|\Psi_{t}\middle|\mathcal{S}_{\nu,a,b}\right\rangle{\Psi_{t}}=1$ (53) indicates nontrivial string order of $\lvert\Psi_{t}\rangle$ (36). ∎

### v3

@@ Appendix I Proof of Lemma 11
where the support of $\Sigma^{\nu}_{1}$ is entirely to the right of $\sigma^{\nu}_{i}$, and thus constitutes an equivalent string order parameter $\mathsf{S}^{\,}_{\nu,t}$ (N.173) of the same form as Eq. 54 with a new left endpoint, where the left-endpoint operators for $\nu=x,z$ are guaranteed to anticommute by Eq. N.172.
Importantly, we observe that the new string-order parameter in Eq. N.173 has the same important properties as the naïve counterpart (54). Both objects have (i) unit expectation value in the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36); (ii) a right-endpoint operator given by $\sigma^{\nu}_{f}$; (iii) a left-endpoint operator of the form $\mathcal{U}\,\tau^{\nu}\,\mathcal{U}^{\dagger}$, where the objects $\tau^{\nu}$ satisfy the Pauli algebra; and (iv) all of the operators between the two endpoint commute for different values of $\nu$.
There are two key features that facilitate iterating this procedure to recover new endpoints. The first is that the left-endpoint operators (N.174) are of the same form as the original Pauli operator in Eq. 54. The second is that there exists a “leftmost” choice of $\Sigma^{\nu}_{1}$ (N.172) such that the clustered observables $\bar{A}^{\,}_{\nu,3}$ through $\bar{A}^{\,}_{\nu,m}$ are not connected to $\Sigma^{\nu}_{1}$ under $\mathcal{U}$. Essentially, a set of equivalent operators recovers upon multiplying $\Sigma^{\nu}_{1}$ (N.172) by stabilizer elements $\mathsf{S}^{\,}_{0}$ of $\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$; the “leftmost” left-endpoint operator (N.174) is the operator $\Sigma^{\nu}_{1}\,\mathsf{S}^{\,}_{0}$ for $\mathsf{S}^{\,}_{0}\in\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$ whose support is closest to—but disjoint from—the initial site $i$ (i.e., whose right weight is closest to $i$).
@@
$\Sigma^{\nu}_{1}\,\Sigma^{\nu}_{2}=\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,2}\,\mathcal{U}\,,~{}~{}$ (N.175)
@@
$\sigma^{\nu}_{i}\,\Sigma^{\nu}_{s}=\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,s}\,\mathcal{U}\,,~{}~{}$ (N.176)
@@
$\Sigma^{\nu}_{s-1}\,\Sigma^{\nu}_{s}=\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,s}\,\mathcal{U}\,,~{}$ (N.177)
@@
One can iterate this procedure until all bulk operators except $\bar{A}^{\,}_{\nu,m}$ have been exhausted. Choosing the leftmost form of each $\Sigma^{\nu}_{s}$ guarantees that each $\Sigma^{\nu}_{s}$ only has support to the right of $\Sigma^{\nu}_{s-1}$, and thus constitutes a genuinely new endpoint. This choice only affects the total number of measurement “regions” $m$, which delineate endpoints.
@@
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t,s}$ $=\left(\mathcal{U}\,\Sigma^{\nu}_{s}\,\mathcal{U}^{\dagger}\right)\,\bar{A}^{\vphantom{\prime}}_{\nu,s+1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\left(\sigma^{\nu}_{f}\right)$ (N.178)
for all $s\in[1,m]$, such that $\left\langle\Psi^{\,}_{t}\middle|\mathsf{S}^{\,}_{\nu,t,s}\middle|\Psi^{\,}_{t}\right\rangle=1$ (60) indicates perfect string order of $\lvert\Psi^{\,}_{t}\rangle$ (36).
We next identify new right endpoints. Returning to Eq. N.171, using the same operators $\Sigma^{\nu}_{s}$ defined in Eqs. N.176 and N.177. Importantly, we take $\Sigma^{\nu}_{s}$ to be the operator $\Sigma^{\nu}_{s}\,\mathsf{S}^{\,}_{0}$ for $\mathsf{S}^{\,}_{0}\in\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$ with the leftmost right weight (i.e., whose support is farthest to the left). All such operators are equivalent, since these quantities are (effectively) evaluated in the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6), so $\mathsf{S}^{\,}_{0}\cong\mathds{1}$.
@@
$\mathsf{S}^{\prime}_{\nu,t,s}$ $\equiv\mathsf{S}^{\vphantom{\prime}}_{\nu,t,0}\,\mathsf{S}^{\dagger}_{\nu,t,s}$
$=\left(\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}\right)$
×$\times\left(\sigma^{\nu}_{f}\,\bar{A}^{\vphantom{\prime}}_{\nu,m}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,s+1}\,\mathcal{U}\,\Sigma^{\nu}_{s}\,\mathcal{U}^{\dagger}\right)$
$=\left(\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\right)\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,s}\,\left(\mathcal{U}\,\Sigma^{\nu}_{s}\,\mathcal{U}^{\dagger}\right)\,\,,~{}~{}$ (N.179)
which has the same left-endpoint operator $\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}$ as $\mathsf{S}^{\,}_{\nu,t,0}$ (54) but distinct right-endpoint operators $\Sigma^{\nu}_{s}$, which are the same endpoint operators in Eq. 57.
@@
$\mathcal{S}^{\vphantom{\prime}}_{\nu,a,b}$ $\equiv\mathsf{S}^{\vphantom{\prime}}_{\nu,t,a}\,\mathsf{S}^{\dagger}_{\nu,t,b}$ (N.180)
$=\left(\mathcal{U}\,\Sigma^{\nu}_{a}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,a+1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}\right)$ (N.181)
×$\times\left(\sigma^{\nu}_{f}\,\bar{A}^{\vphantom{\prime}}_{\nu,m}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,b+1}\,\mathcal{U}\,\Sigma^{\nu}_{b}\,\mathcal{U}^{\dagger}\right)$ (N.182)
$=\left(\mathcal{U}\,\Sigma^{\nu}_{a}\,\mathcal{U}^{\dagger}\right)\,\bar{A}^{\vphantom{\prime}}_{\nu,a+1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,b}\,\left(\mathcal{U}\,\Sigma^{\nu}_{b}\,\mathcal{U}^{\dagger}\right)\,\,,~{}~{}$ (57)
and since the support of $\Sigma^{\nu}_{s+1}$ is entirely to the right of $\Sigma^{\nu}_{s}$, the support of $\Sigma^{\nu}_{s}$ is to the left of $\Sigma^{\nu}_{s+1}$. Hence, the operators $\Sigma^{\nu}_{s}$ constitute distinct left and right endpoint operators for each $s\in[1,m]$, where $\left\langle*\middle|\Psi^{\,}_{t}\middle|\mathcal{S}^{\,}_{\nu,a,b}\right\rangle{\Psi^{\,}_{t}}=1$ (53) indicates nontrivial string order of $\lvert\Psi^{\,}_{t}\rangle$ (36). ∎